Make a sketch of the Gibbs free energy, G, versus temperature, T, for a single component system as the temperature increases at a fixed pressure, P.

Indicate the temperatures where the pure substance changes from liquid to solid, and from solid to gas, and comment on the slopes of the plot of G versus T.
Answer #1


Solid Liquid Gas Solid Liquid Gas stablestable stable T. Temperature, TdH dT 298 K TST dG Temperaturę TG=H-TS dG = dH-TdS-SdT dG = dE + PdV +VdP-Tds-sal. since H = E+ PV sınce Be OG So at a constant pressure the slope of the free energy curve